Misty has $2.35 in dimes and nickles. She has one less nickel thsn dimes. How many of each coin does she have?
Answer by stanbon(57214)

Misty has $2.35 in dimes and nickles. She has one less nickel then dimes. How many of each coin does she have?
-----------------
Let # of dimes be "x"; value of these is 10x cents
-----------
# of nickels is "x-1"; value of these is 5(x-1)=5x-5 cents
-----------
EQUATION:
value + value = 235 cents
10x + 5x-5 = 235
15x = 240
x = 16 (# of dimes)
x-1=15 (# of nickels)
================
Cheers,
Stan H.